### v1

# Bogomolov property for certain infinite non-Galois extensions

Abstract.
For an algebraic number $\alpha$, let $h(\alpha)$ denote its logarithmic Weil height. In 2002, Bombieri and Zannier obtained lower bounds for Weil height for totally $p$-adic extensions of $\mathbb{Q}$, i.e., infinite Galois extension over $\mathbb{Q}$ with finite local degree over a rational prime $p$. In this paper, we formulate and prove the non-Galois analogue of this result. In particular, if a positive proportion of conjugates of $\alpha\in\overline{\mathbb{Q}}$ lie in a finite local extension $K_{\nu}/\mathbb{Q}_{p}$, then we obtain a lower bound for $h(\alpha)$. We also discuss the lower bound for canonical height of an elliptic curve $E/\mathbb{Q}$ in such infinite extensions.
Key words and phrases:
Weil height, Bogomolov property, Northcott property, Lehmer’s conjecture, Canonical height of elliptic curves
2020 Mathematics Subject Classification:
11G50, 11M41, 11R04, 11R06
1. Introduction
Let $K/\mathbb{Q}$ be a number field. For $\alpha\in K^{*}$, the absolute logarithmic height or logarithmic Weil height is defined as
$h(\alpha)=\sum_{v\in M_{K}}\log^{+}|\alpha|_{v},$
where $M_{K}$ is the set of places of $K$, $\log^{+}x=\max(0,\log x)$ and $|\alpha|_{v}$ is the normalized valuation on $\alpha$ defined as follows. For non-archimedean valuation $v$,
$|\alpha|_{v}:=\left(N\mathfrak{p}\right)^{-\frac{ord_{\mathfrak{p}}(\alpha)}{[K\,:\,\mathbb{Q}]}},$
where $\mathfrak{p}$ is the prime ideal corresponding to the valuation $v$. If $v$ is archimedean, then
$|\alpha|_{v}:=|\sigma(\alpha)|^{\frac{1}{[K\,:\,\mathbb{Q}]}},$
where $\sigma$ is the embedding of $\alpha$ corresponding to $v$.
Kronecker’s well-known theorem tells us that an algebraic number $\alpha$ satisfies $h(\alpha)=0$ if and only if $\alpha$ is a root of unity. In 1933, Lehmer [19] conjectured that if a non-zero $\alpha\in\overline{\mathbb{Q}}$ is not a root of unity, then
$h(\alpha)\geq\frac{c}{[\mathbb{Q}(\alpha):\mathbb{Q}]}$
for an absolute constant $c>0$. Lehmer’s conjecture can also be expressed in terms of the Mahler measure of $\alpha$. Let $f(x)=a_{n}x^{n}+\cdots+a_{1}x+a_{0}\in\mathbb{Z}[x]$ be the minimal polynomial of $\alpha$. Then, its Mahler measure is defined as
$M(\alpha):=|a_{n}|\prod_{i}\max(1,|\alpha_{i}|),$
where $\alpha_{i}$’s denote the conjugates of $\alpha$. Mahler measure is connected to the height of $\alpha$ by the relation
$\log M(\alpha)=h(\alpha)\,[\mathbb{Q}(\alpha):\mathbb{Q}].$
Thus, Lehmer’s conjecture can be stated as
$\log M(\alpha)\geq c$
for non-zero $\alpha\in\overline{\mathbb{Q}}$ not a root of unity and $c>0$ an absolute constant.
While this conjecture remains unresolved, there has been significant progress in recent times. The interested reader may refer to the excellent survey articles [31], [33] and the book by D. Masser [21] for a comprehensive account of this problem.
A well-known theorem of Northcott [25] states that there are only finitely many algebraic numbers, not roots of unity, with bounded degree and height. This inspires the definition of Northcott property (N). A set $S\subseteq\overline{\mathbb{Q}}$ is said to satisfy the Northcott property (N) if for any $c>0$,
$\{\alpha\in S\,|\,\alpha\text{ non-zero and not a root of unity}:h(\alpha)<c\}$
is a finite set. Another related property on a set of algebraic numbers is the Bogomolov property (B). We say a set $S\subset\overline{\mathbb{Q}}$ satisfies the Bogomolov property (B) if there exists a constant $c>0$, such that
$\{\alpha\in S\,|\,\alpha\text{ non-zero and not a root of unity}:h(\alpha)<c\}$
is an empty set. If a set $S$ satisfies property (N), then it clearly satisfies property (B). Equivalently, $S$ has Northcott (resp. Bogomolov) if
$\liminf_{\alpha\in S}h(\alpha)=\infty\,\,\,\,\left(\text{resp}.\,\liminf_{\alpha\in S}h(\alpha)>0\right).$
Moreover, if $S$ satisfies property (B), then there is an absolute lower bound for $h(\alpha)$ with $\alpha\in S$ not a root of unity, and hence Lehmer’s conjecture holds for $S$. For instance, F. Amoroso and R. Dvornicich [1] have proved that $\mathbb{Q}^{ab}$, the maximal abelian extension over $\mathbb{Q}$ satisfies property (B). They showed that for non-zero $\alpha\in\mathbb{Q}^{ab}$, not a root of unity,
$h(\alpha)>\frac{\log 5}{12}.$
This was generalized to $K^{ab}$ by F. Amoroso and U. Zannier [3]. Earlier, in 1973, Schinzel [26] obtained property (B) for the set $\mathbb{Q}^{tr}$ of totally real numbers. More precisely, for $\alpha\in\mathbb{Q}^{tr}\setminus\{\pm 1\}$, he proved that
$h(\alpha)\geq\frac{1}{2}\log\left(\frac{1+\sqrt{5}}{2}\right).$
In fact, the angular equidistribution theorem by Amoroso-Mignotte [2]( also see [21, Theorem 15.2], M. Mignotte [23] and Y. Bilu[7]) implies that all $\alpha\in\overline{\mathbb{Q}}$, whose conjugates are not equidistributed in a given sector on the complex plane satisfy property (B). In other words, as long as the conjugates do not have angular equidistribution, we have property (B). A natural question to ponder is whether one can formulate a non-archimedean counterpart to this equidistribution property. In this paper, we formulate and prove such an equidistribution theorem.
In 2001, E. Bombieri and U. Zannier [8] obtained Bogomolov property for infinite extensions with finite local degree, also called totally $p$-adic fields of type $(e,f)$. Let $\mathcal{L}$ over $\mathbb{Q}$ be an infinite Galois extension. Define
$S(\mathcal{L}):=\{p\,\text{ prime}\,:\,[\mathcal{L}_{v}:\mathbb{Q}_{p}]<\infty\,\,\text{for some}\,\,v\,\,\text{above}\,\,p\}.$
Then, they showed that
$\liminf_{\alpha\in\mathcal{L}}\,\,h(\alpha)\,\,\geq\frac{1}{2}\sum_{p\in S(\mathcal{L})}\frac{\log p}{e_{p}\,\left(p^{f_{p}}+1\right)},$ (1)
where $e_{p}$ is the ramification index and $f_{p}$ is the residual degree of $\mathcal{L}_{v}/\mathbb{Q}_{p}$. Thus, if $S(\mathcal{L})$ is non-empty, then $\mathcal{L}$ satisfies property (B). Furthermore, if the right-hand side in (1) diverges, then $\mathcal{L}$ satisfies property (N). Note that if $p\in S(\mathcal{L})$, then for any $\alpha\in\mathcal{L}$, all its conjugates lie in some finite extension $L_{\nu}$ of $\mathbb{Q}_{p}$. This could be thought of as the analogue of all the roots being real in the archimedean case. Naturally, a question of interest is to understand what happens when a positive proportion of roots lie in $L_{\nu}$, a finite extension of $\mathbb{Q}_{p}$.
There have been several refinements of (1) in recent times. In 2015, P. Fili and C. Petsche [12], used methods of potential theory to obtain lower bounds on the height of algebraic numbers, all whose conjugates lie in certain local fields. Later, these methods were extended by P. Fili and I. Pritsker [13] to obtain several new lower bounds for algebraic numbers all of whose conjugates lie in real intervals and p-adic discs. Despite this, the question posed in the previous paragraph is unknown.
In order to state our main theorems, we introduce the following notation. For a number field $K/\mathbb{Q}$ and a rational prime power $q$, define
$\mathcal{N}_{q}(K):=\text{ the number of prime ideals of }K\text{ with norm }q.$
An infinite extension $\mathcal{L}/\mathbb{Q}$ can be written as a tower of number fields
$\mathcal{L}\supsetneq\cdots\supsetneq L_{m}\supsetneq L_{m-1}\supsetneq\cdots\supsetneq L_{1}=\mathbb{Q},$
where $L_{i}/\mathbb{Q}$ are finite extensions. For any rational prime power $q=p^{k}$, define
$\psi_{q}(\mathcal{L})=\psi_{q}:=\lim_{i\to\infty}\frac{\mathcal{N}_{q}(L_{i})}{[L_{i}:\mathbb{Q}]},$
where $\mathcal{N}_{q}(L)$ is the number of prime ideals in $L$ with norm $q$. This limit is well-defined (as shown in Proposition 4.2), i.e., independent of the tower $\{L_{i}\}$ and $0\leq\psi_{q}\leq 1$. For instance, if a prime $p$ splits completely in $\mathcal{L}$, then $\psi_{p}=1$ and $\psi_{p^{k}}=0$ for all $k>1$. If $\psi_{q}>0$ for some prime power $q$, we call the extension $\mathcal{L}$ to be asymptotically positive. This setup is inspired by the theory of asymptotically exact families by M. A. Tsfasman, S. G. Vlăduţ [32]. Our main theorem is to prove the Bogomolov property (B) for asymptotically positive extensions.
Theorem 1.1.
Let $\mathcal{L}/\mathbb{Q}$ be an infinite extension of $\mathbb{Q}$. Then,
$\liminf_{\alpha\in\mathcal{K}}h(\alpha)\geq\frac{1}{2}\sum_{q}\psi_{q}\,\frac{\log q}{q+1},$
where $q$ runs over all prime powers. Therefore, if $\mathcal{L}$ is asymptotically positive, then $\mathcal{L}$ has property (B).
The above theorem settles Conjecture 2.2 from [10], which was proposed by the authors earlier, in an attempt to understand (1) for non-Galois extensions. For an infinite Galois extension $\mathcal{L}/\mathbb{Q}$, if $p\in S(\mathcal{L})$ with ramification index $e_{p}$ and residue class degree $f_{p}$, then
$\frac{\log p}{e_{p}(p^{f_{p}}+1)}=\frac{\log p^{f_{p}}}{e_{p}f_{p}(p^{f_{p}}+1)}=\psi_{p^{f_{p}}}\frac{\log p^{f_{p}}}{p^{f}+1}.$
Thus, specializing to Galois extensions, Theorem 1.1 is equivalent to (1). Note that for $\mathcal{L}/\mathbb{Q}$ Galois, the condition $[\mathcal{L}_{v}:\mathbb{Q}_{p}]<\infty$ is equivalent to saying that there exists a finite Galois extension $L/\mathbb{Q}$ with $L\subset\mathcal{L}$ such that all the prime ideals above $p$ in $L$ split completely in $\mathcal{L}$. However, it is difficult to obtain similar results for non-Galois extensions, where we do not have eventual complete splitting. Naively taking the Galois closure and applying (1) often leads to poor bounds on $h(\alpha)$. We illustrate this with an example below.
Example. Fix a rational prime $p$. We construct an infinite extension $\mathcal{L}/\mathbb{Q}$ such that $\psi_{p}>0$, but $p\notin S(\widetilde{\mathcal{L}})$, where $\widetilde{\mathcal{L}}$ is the Galois closure of $\mathcal{L}$ over $\mathbb{Q}$.
Let $K/\mathbb{Q}$ be a number field with distinct prime ideals $P_{1},P_{2},...,P_{s}$ and $Q_{1},Q_{2},...,Q_{t}$. For any integer $n\geq 2$, there exists a degree $n$ extension $L$ of $K$ such that $P_{1},P_{2},...,P_{s}$ split completely and $Q_{1},Q_{2},...,Q_{t}$ totally ramify over $L$. Let $L_{1}/\mathbb{Q}$ be a degree $4$ extension where $p$ splits completely into ${\mathfrak{p}}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3},\mathfrak{p}_{4}$. Construct $L_{2}/L_{1}$ a degree $9$ extension such that $\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3}$ split completely and $\mathfrak{p}_{4}$ ramifies, but not totally in $L_{2}$. This is achieved by first constructing $L_{2}^{\prime}/L_{1}$ of degree $3$ such that $\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3}$ split completely and $\mathfrak{p}_{4}$ ramifies completely. Then, we construct $L_{2}/L_{2}^{\prime}$ where all primes in $L_{2}^{\prime}$ above $\mathfrak{p}_{1},\mathfrak{p}_{2}$ and $\mathfrak{p}_{3}$ split completely and the prime above $\mathfrak{p}_{4}$ also splits completely. Inductively construct $L_{i+1}/L_{i}$, an extension of degree $p_{i+1}^{2}$, where $p_{i}$ is the $i$-th prime number, such that $\mathcal{N}_{p}(L_{i})-\frac{\mathcal{N}_{p}(L_{i})}{{p_{i}}^{2}}$ of the prime ideals with norm $p$ split completely in $L_{i+1}$ and the remaining prime ideals with norm $p$ ramify, but not completely in $L_{i+1}$. For this tower =$\mathcal{L}=\{L_{i}\}$, note that
$\psi_{p}=\lim_{i\to\infty}\frac{\mathcal{N}_{p}(L_{i})}{[L_{i}:\mathbb{Q}]}\geq\prod_{p_{i}}\left(1-\frac{1}{{p_{i}}^{2}}\right)=\frac{1}{\zeta(2)},$
where the product above runs over all rational primes $p_{i}$. Thus, we have $\psi_{p}>0$. However, since the ramification degree of $p$ is unbounded, $p\notin S(\widetilde{\mathcal{L}})$.
In special cases, where there is minimal ramification of a prime in the infinite extension, it is feasible to obtain finer lower bounds, which could potentially surpass Theorem 1.1. An infinite extension $\mathcal{L}/\mathbb{Q}$ is said to be almost split at a prime $p$ if $\psi_{p}=1$ and almost unramified at a prime $p$ if
$\sum_{\begin{subarray}{c}q=p^{k}\\ k\geq 1\end{subarray}}k\,\psi_{q}=1.$
As an immediate consequence of Theorem 1.1, we have
Corollary 1.2.
Let $\mathcal{L}/\mathbb{Q}$ be an infinite extension which is almost split at a prime $p$. Then
$\liminf_{\alpha\in\mathcal{L}}h(\alpha)\geq\frac{\log p}{2(p+1)}.$
We also establish a refined lower bound for almost unramified extensions.
Theorem 1.3.
Let $\mathcal{L}/\mathbb{Q}$ be an infinite extension which is almost unramified at a prime $p$. Then
$\liminf_{\alpha\in\mathcal{L}}h(\alpha)\geq\sum_{\begin{subarray}{c}q=p^{k}\\ k\geq 1\end{subarray}}k^{2}\,{\psi_{q}}^{2}\frac{\log q}{2(q+1)}.$
2. Almost splitting at a prime $p$
For ease of exposition, we first prove Corollary 1.2. The proof of Theorem 1.1 requires the development of the theory of relative heights, which is done in Section 3.
This proof of Corollary 1.2 is inspired by the proof of (1) due to Bombieri-Zannier in [8]. For a number field $L/\mathbb{Q}$, denote by $n_{L}$, the degree $[L:\mathbb{Q}]$. Let $\mathcal{L}$ be an infinite extension which is almost split at prime $p$, i.e., $\psi_{p}=1$. Then, for $\alpha\in\mathcal{L}$ and $L=\mathbb{Q}(\alpha)$,
$\mathcal{N}_{p}(L)=n_{L}+o(n_{L})$
as $n_{L}$ tends to infinity. We start by showing that the number of ramified primes over $p$ is small.
Lemma 2.1.
Let $\mathcal{L}$ be almost split at $p$. For any number field $L\subseteq\mathcal{L}$, the number of prime ideals in $L$ which ramify over $p$ is $o(n_{L})$ as $n_{L}$ tends to infinity.
Proof.
Let $\{\mathfrak{p}_{1},\mathfrak{p}_{2},\cdots,\mathfrak{p}_{j}\}$ denote the prime ideals in $L$ above $p$. Recall that $n_{L}=\sum_{j}e_{j}f_{j}$, where $e_{j}$ is the ramification index and $f_{j}$ is the residue class degree of $\mathfrak{p}_{j}$ over $p$. Define
$A:=\{\mathfrak{p}\in L\,:\,\mathfrak{p}\mid p\mathcal{O}_{L},Norm(\mathfrak{p})=p,e(\mathfrak{p})=1\},$
and
$B:=\{\mathfrak{p}\in L\,:\,\mathfrak{p}\mid p\mathcal{O}_{L},Norm(\mathfrak{p})=p,e(\mathfrak{p})\geq 2\}.$
Since $\mathcal{L}$ is almost split at $p$ and $L\subset\mathcal{L}$, we have $\mathcal{N}_{p}(L)=|A|+|B|=n_{L}+o(n_{L})$. Also $f(\mathfrak{p})=1$ for all prime ideals $\mathfrak{p}\in A\cup B$. Hence,
$n_{L}=\sum_{j}e_{j}f_{j}=|A|+|B|+\sum_{\mathfrak{p}_{j}\in B}(e_{j}-1)|B|+\sum_{\mathfrak{p}_{l}\notin A\cup B}e_{l}f_{l}.$
Thus, we conclude that $|B|=o(n_{L})$ and $|A|=n_{L}+o(n_{L})$ as $n_{L}$ tends to infinity. ∎
We now recall the classical Mahler’s inequality [20].
Theorem 2.2 (Mahler).
Let $f(x)=a_{n}x^{n}+\cdots+a_{1}x+a_{0}\in\mathbb{C}[x]$ be a polynomial with roots $\alpha_{1},\alpha_{2},\cdots,\alpha_{n}$. Let
$D:=a_{n}^{2n-2}\,\prod_{i>j}(\alpha_{i}-\alpha_{j})^{2}$
be its discriminant. Then,
$|D|\leq n^{n}M(f)^{2n-2},$
where $M(f)$ denotes the Mahler measure of $f$, given by $M(f):=|a_{n}|\prod_{i}\max(1,|\alpha_{i}|)$.
We also recall the following well known lemma from algebraic number theory (see [24, Prop. II. 8.2])
Lemma 2.3.
Let $K$ be a number field. Suppose $L=K(\alpha)$ and $f_{K}(x)\in\mathcal{O}_{K}[x]$ be the minimal polynomial of $\alpha$ over $K$. For a prime ideal $\mathfrak{p}\in K$, if $\mathfrak{p}\mathcal{O}_{L}=\mathfrak{q}_{1}^{e_{1}}\mathfrak{q}_{2}^{e_{2}}\cdots\mathfrak{q}_{g}^{e_{g}}$ and $f_{K}(x)$ factors in $K_{\nu}$, the local field corresponding to $\mathfrak{p}$, as $f_{K}(x)=f_{1}(x)f_{2}(x)\cdots f_{G}(x)$, then $g=G$ and upto ordering $\deg f_{j}(x)=e_{j}f_{j}$, where $e_{j}$ and $f_{j}$ are the ramification index and the residue class degree of $\mathfrak{q}_{j}$ respectively.
In particular, let $L=\mathbb{Q}(\alpha)$ and $f(x)\in\mathbb{Z}[x]$ be the minimal polynomial of $\alpha$. If $p\mathcal{O}_{L}=\mathfrak{p}_{1}^{e_{1}}\mathfrak{p}_{2}^{e_{2}}\cdots\mathfrak{p}_{g}^{e_{g}}$ and $f(x)$ factors in $\mathbb{Q}_{p}$ as $f(x)=f_{1}(x)f_{2}(x)\cdots f_{G}(x)$, then $g=G$ and upto ordering $\deg f_{j}(x)=e_{j}f_{j}$.
Proof of Corollary 1.2.
Let $\mathcal{L}$ be almost split at a prime $p$. By Northcott’s theorem, there are finitely many $\alpha$ with bounded degree such that $h(\alpha)<\frac{\log p}{2(p+1)}$. For any $\alpha\in\mathcal{L}$, let $L=\mathbb{Q}(\alpha)$ and $f(x)\in\mathbb{Z}[x]$ be its minimal polynomial of degree $n$. By Lemma 2.1 and 2.3, the minimal polynomial $f$ factorizes over $\mathbb{Q}_{p}$ as
$f(x)=a_{n}(x-\alpha_{1})(x-\alpha_{2})\cdots(x-\alpha_{t})g(x),$
where $t=n+o(n)$ and $\deg g(x)=o(n)$ as $n$ goes to infinity. Thus the $p$-adic valuation $\nu$ on $\mathbb{Q}_{p}$ can be applied to $\alpha_{1},\cdots,\alpha_{t}$. Let $L_{\omega}$ be the splitting field of $f(x)$ over $\mathbb{Q}_{p}$ with valuation $\omega$. Write
$f(x)=a_{n}(x-\beta_{1})(x-\beta_{2})\cdots(x-\beta_{n}),$
where $\beta_{i}\in L_{\omega}$ satisfying
$\omega(\beta_{1})\geq\cdots\geq\omega(\beta_{r})\geq 0\geq\omega(\beta_{r+1})\geq\cdots\geq\omega(\beta_{n}).$
Note that except for $o(n)$ many $\beta_{i}$’s, the rest are same as $\alpha_{i}$’s up to ordering. Consider the discriminant
$D(f)$ $=a_{n}^{2n-2}\prod_{i<j}(\beta_{i}-\beta_{j})^{2}.$
The contribution in the product where at least one $\omega(\beta_{j})<0$ can be bounded by
$\omega\left(\prod_{r+1}^{n}\prod_{i=1}^{j-1}(\beta_{j}-\beta_{i})\right)\geq\sum_{j=r+1}^{n}(j-1)\omega(\beta_{j}).$
Hence, we obtain
$\omega(D(f))$ $\geq(2n-2)\omega(a_{n})+2\sum_{0<i<j\leq r}\omega(\beta_{j}-\beta_{i})+2\sum_{j=r+1}^{n}(j-1)\omega(\beta_{j})$
$\geq 2\sum_{i<j\leq r}\omega(\beta_{j}-\beta_{i})-2\sum_{j=r+1}^{n}(m-j)\omega(\beta_{j}).$ (2)
Recall that $f(x)$ has $n+o(n)$ roots in $\mathbb{Q}_{p}$. Let $e_{\omega}$ be the ramification index of $L_{\omega}/\mathbb{Q}_{p}$. For $\beta\in\mathbb{Q}_{p}$, we have $\omega(\beta)=\nu(\beta)\,e_{\omega}$. Since all terms in the RHS of (2) are positive, we can write
$\nu(D(f))$ $=\frac{\omega(D(f))}{e_{\omega}}\geq\frac{2}{e_{\omega}}\sum_{i<j\leq r}\omega(\beta_{j}-\beta_{i})-\frac{2}{e_{\omega}}\sum_{j=r+1}^{n}(m-j)\omega(\beta_{j})$
$\geq 2\sum_{\begin{subarray}{c}i<j\leq r\\ \beta_{i},\beta_{j}\in\mathbb{Q}_{p}\end{subarray}}\nu(\beta_{j}-\beta_{i})-2\sum_{\begin{subarray}{c}r<j\leq n\\ \beta_{j}\in\mathbb{Q}_{p}\end{subarray}}(n-j)\nu(\beta_{j}).$ (3)
Suppose $N_{x}$ denote the number of roots $\beta_{j}$ of $f(x)$ in $\mathbb{Q}_{p}$, which lie in the residue class $x\bmod p$. If $\beta_{i},\beta_{j}\in\mathbb{Q}_{p}$ lie in the same residue class modulo $p$, then $\nu(\beta_{i}-\beta_{j})\geq 1$. Hence,
$\sum_{\begin{subarray}{c}i<j\leq r\\ \beta_{i},\beta_{j}\in\mathbb{Q}_{p}\end{subarray}}\nu(\beta_{j}-\beta_{i})\geq\sum_{x\in\mathbb{F}_{p}}\frac{N_{x}(N_{x}-1)}{2}.$
Therefore, by (2), we have the lower bound
$\nu(D(f))\geq\sum_{x\in\mathbb{F}_{p}}N_{x}(N_{x}-1)+(n-r)(n-r-1).$
Since the number of roots of $f(x)$ in $\mathbb{Q}_{p}$ is nothing but $n-r+\sum_{x\in\mathbb{F}_{p}}N_{x}=n+o(n)$, applying Cauchy-Schwarz inequality, we obtain
$\nu(D(f))$ $\geq\sum_{x\in\mathbb{F}_{p}}N_{x}(N_{x}-1)+(n-r)(n-r+1)$
$=\sum_{x\in\mathbb{F}_{p}}N_{x}^{2}+(n-r)^{2}-n+o(n)$
$\geq\frac{(n+o(n))^{2}}{p+1}+O(n)$
$=\frac{n^{2}}{p+1}+o(n^{2}).$
Therefore,
$\log D(f)\geq n^{2}\,\left(\frac{\log p}{p+1}\right)+o\left(n^{2}\frac{\log p}{p}\right).$
Applying Theorem 2.2, we obtain
$h(\alpha)=\frac{\log M(\alpha)}{n}$ $\geq\frac{(n-1)\log M(f)}{n^{2}}\geq\frac{\log D(f)}{2n^{2}}-\frac{\log n}{n}$
$\geq\frac{\log p}{2(p+1)}+o\left(\frac{\log p}{p}\right)-\frac{\log n}{n}.$
As $n\to\infty$, we have
$\liminf_{\alpha\in\mathcal{L}}h(\alpha)\geq\frac{\log p}{2(p+1)}$
as required. This proves Corollary 1.2.
3. Relative Weil height and Mahler’s inequality
In this section, we develop the theory of relative Weil height and prove an analog of Mahler’s inequality (Theorem 2.2), which is essential for the proof of Theorem 1.1 and 1.3.
Let $K/\mathbb{Q}$ be a number field and $\alpha\in\overline{\mathbb{Q}}\,\setminus\,K$ be an algebraic number of degree $n=[\mathbb{Q}(\alpha):\mathbb{Q}]$. Let
$f(x)=a_{n}x^{n}+a_{n-1}x^{n-1}+\ldots+a_{0}$
be the minimal polynomial of $\alpha$ over $\mathbb{Q}$ and $\alpha_{1},\alpha_{2},\ldots,\alpha_{n}$ be the distinct complex roots of $f$. Recall that the Mahler measure and the logarithmic Weil height of $\alpha$ is given by
$M(\alpha)=|a_{n}|\prod_{i=1}^{n}\operatorname{max}\{1,|\alpha_{i}|\}\qquad\text{ and }\qquad h(\alpha)=\frac{\log M(\alpha)}{n}.$
Let $f_{K}(x)=b_{m}x^{m}+b_{m-1}x^{m-1}+\ldots+b_{0}$ be the minimal polynomial of $\alpha$ over $K$ of degree $m=[K(\alpha):K]$ and $\alpha_{j}$’s be its roots in the algebraic closure of $K$. We define the relative Mahler measure of $\alpha$ with respect to an embedding $\sigma:K\hookrightarrow\mathbb{C}$ as
$M_{\sigma(K)}(\alpha)=|\sigma(b_{m})|\prod_{i=1}^{m}\operatorname{max}\{1,|\alpha_{j}^{\sigma}|\},$
where $\{\alpha_{j}^{\sigma}\}_{j=1}^{m}\,s$ are complex roots of $\sigma(f_{K}(x))$. Counting all the roots with multiplicity, one obtains
${f_{K}(x)}^{[K(\alpha):\mathbb{Q}(\alpha)]}=\prod_{\sigma}\sigma(f_{K}(x)),$ (4)
where $\sigma$ runs over all the complex embeddings of $K$.
$\mathbb{Q}$$K$$K(\alpha)$$\mathbb{Q}(\alpha)$nm
Thus a natural way to define the relative height of $\alpha$ with respect to an embedding $\sigma:K\hookrightarrow\mathbb{C}$ is
$h_{\sigma(K)}(\alpha)=\frac{\log M_{\sigma(K)}(\alpha)}{[K(\alpha):K]}=\frac{\log M_{\sigma(K)}(\alpha)}{m}.$
Using the multiplicativity of Mahler measure, from (4) we have
$h(\alpha)=\frac{1}{[K:\mathbb{Q}]}\sum_{\sigma}h_{\sigma(K)}(\alpha).$ (5)
In other words, the Weil height is the average of relative heights with respect to all the embeddings.
3.1. Mahler’s inequality for relative height
Let $f_{K}(x)=a_{m}\,x^{m}+a_{m-1}\,x^{m-1}+\cdots+a_{0}$ be the minimal polynomial of $\alpha$ over $K$ and $\alpha_{1},\alpha_{2},\ldots,\alpha_{m}$ be its conjugate over an algebraic closure $\overline{K}$. Then, the discriminant of $f_{K}$ is given as
$D(f_{K})={a_{m}}^{2m-2}\,\prod_{i<j\leq m}\,\left(\alpha_{i}-\alpha_{j}\right)^{2}.$
Thus for an embedding $\sigma$ of $K$ in $\mathbb{C}$, we have
$|\sigma(D(f_{K}))|=|\sigma({a_{m}})|^{2m-2}\,\,\prod_{i<j\leq m}\,|\sigma(\alpha_{i})-\sigma(\alpha_{j})|^{2}.$
We now proceed as in the proof of Mahler’s inequality [20]. The product in the last equality is square of absolute value of a Vandermonde determinant. Thus applying Hadamard’s inequality, we obtain
$|\sigma(D(f_{K}))|$ $\leq|\sigma({a_{m}})|^{2m-2}\,\,\prod_{i=1}^{m}\,\left(\,\sum_{j=0}^{m-1}|{\alpha_{i}}^{j}|^{2}\,\right)$
$\leq|\sigma({a_{m}})|^{2m-2}\,\,\prod_{i=1}^{m}\,m\,\operatorname{max}\{1,|\alpha_{i}|^{2m-2}\}$
$=m^{m}M_{\sigma(K)}(\alpha).$
Taking logarithm, we deduce Mahler’s inequality for relative height.
Lemma 3.1.
Let $\alpha\in\overline{\mathbb{Q}}\setminus K$ be an algebraic number, with minimal polynomial $f_{K}(x)$ over $K$ with degree $m$. Then, for any complex embedding $\sigma$ of $K$, we have
$\log|\sigma(D(f_{K})|\,\,\leq\,\,m\log m+(2m-2)\log M_{\sigma(K)}(\alpha).$
In other words, for relative Weil height
$h_{\sigma(K)}(\alpha)\geq\frac{1}{2}\left(\frac{\log|\sigma(D(f_{K})|}{m^{2}}-\frac{\log m}{m}\right).$
Summing over all complex embeddings $\sigma$ of $K$ and using (5), we deduce that
$h(\alpha)\geq\frac{1}{2\,[K:\mathbb{Q}]}\left(\frac{\log\left(|N_{K/\mathbb{Q}}\left(D(f_{K})\right)|\right)}{m^{2}}\right)-\frac{\log m}{m}\,.$ (6)
Remark. It is often advantageous to work with the norm of the relative discriminant over a number field $K$, as opposed to considering the discriminant over $\mathbb{Q}$. This is mainly because we have the identity
$D(f)=|disc(K/\mathbb{Q})|^{2}N_{K/\mathbb{Q}}(D(f_{K})),$
where $f$ and $f_{K}$ are the minimal polynomials of an algebraic number $\alpha$ over $\mathbb{Q}$ and $K$ respectively and $disc(K/\mathbb{Q})$ is the discriminant of $K/\mathbb{Q}$. Thus $|N_{K/\mathbb{Q}}(D(f_{K}))|$ is smaller than $|D(f)|$ and hence, it is more feasible to obtain upper bounds for this norm.
4. Asymptotically positive towers
In this section we state and prove some basic properties of asymptotically positive extensions and the invariant $\psi_{q}$.
Lemma 4.1.
Let $L/K$ be an extension of number fields. Then,
$\sum_{q\leq x}\frac{\mathcal{N}_{q}(K)\,\log q}{[K:\mathbb{Q}]}\geq\sum_{q\leq x}\frac{\mathcal{N}_{q}(L)\,\log q}{[L:\mathbb{Q}]},$
where $q$ runs over all prime powers.
Proof.
If a prime ideal $\mathfrak{q}$ in $L$ has norm $\leq x$, then the prime ideal below, $\mathfrak{p}=\mathfrak{q}\cap K$ also has norm $\leq x$. Suppose a prime ideal $\mathfrak{p}$ in $K$ splits into $\{\mathfrak{q}_{1},\mathfrak{q}_{2},\cdots,\mathfrak{q}_{r}\}$ in $L$. Then,
$\prod_{i}N(\mathfrak{q}_{i})=N(\mathfrak{p})^{[L\,:\,K]}.$
Therefore,
$\sum_{m=1}^{n}m\,\,\mathcal{N}_{p^{m}}(L)\leq[L:K]\sum_{m=1}^{n}m\,\,\mathcal{N}_{p^{m}}(K).$
Mutliplying by $\log p$ and dividing by $[L:\mathbb{Q}]$, we obtain for any prime $p$
$\sum_{q\leq x,q=p^{k}}\frac{\mathcal{N}_{q}(K)\,\log q}{[K:\mathbb{Q}]}\geq\sum_{q\leq x,q=p^{k}}\frac{\mathcal{N}_{q}(L)\,\log q}{[L:\mathbb{Q}]}.$
Summing over all primes $p\leq x$, we obtain the lemma.
Hence, we deduce that for any tower =$\mathcal{L}=\{L_{i}\}$ and any $x>1$, the limit
$\lim_{i\to\infty}\sum_{q\leq x,q=p^{k}}\frac{\mathcal{N}_{q}(L_{i})\,\log q}{[L_{i}:\mathbb{Q}]}$
exists. Therefore, inductively, we can conclude that the limit
$\psi_{q}=\lim_{i\to\infty}\frac{\mathcal{N}_{q}(L_{i})}{[L_{i}:\mathbb{Q}]},$
exists for all prime powers $q$ and takes value between 00 and $1$. We now show that this invariant is independent of the choice of the tower and hence is well-defined.
Proposition 4.2.
Let $\mathcal{L}/\mathbb{Q}$ be an infinite extension and $q$ be a rational prime power. Suppose $\{L_{i}\}$ and $\{K_{i}\}$ are two towers of number fields such that $\mathcal{L}=\bigcup_{i}K_{i}=\bigcup_{i}L_{i}$. Then,
$\lim_{i\to\infty}\frac{\mathcal{N}_{q}(L_{i})}{n_{L_{i}}}=\lim_{i\to\infty}\frac{\mathcal{N}_{q}(K_{i})}{n_{K_{i}}}.$
Proof.
Let $\psi_{q}:=\lim_{i\to\infty}\frac{\mathcal{N}_{q}(K_{i})}{n_{K_{i}}}$ and $\psi_{q}^{\prime}:=\lim_{i\to\infty}\frac{\mathcal{N}_{q}(L_{i})}{n_{L_{i}}}$. Suppose $K_{i}=\mathbb{Q}(\alpha)$. Since $K_{i}\subset\mathcal{L}=\bigcup_{j}L_{j}$, there exists $L_{j}$ such that $\alpha\in L_{j}$. Hence, $K_{i}\subset L_{j}$ for some $j$. Similarly, every $L_{j}\subset K_{l}$ for some $l$. Therefore, applying Lemma 4.1, for every $i$, there exists $j$ and $l$ such that using
$\sum_{q\leq x,q=p^{k}}\frac{\mathcal{N}_{q}(K_{i})\,\log q}{[K_{i}:\mathbb{Q}]}\geq\sum_{q\leq x,q=p^{k}}\frac{\mathcal{N}_{q}(L_{j})\,\log q}{[L_{j}:\mathbb{Q}]}\geq\sum_{q\leq x,q=p^{k}}\frac{\mathcal{N}_{q}(K_{l})\,\log q}{[K_{l}:\mathbb{Q}]}.$
Since $\psi_{q}$ and $\psi_{q}^{\prime}$ exist for all prime powers $q$, taking $i$ and $l$ to infinity, we conclude that for all $x>1$,
$\sum_{q\leq x}\psi_{q}\log q\geq\sum_{q\leq x}\psi_{q}^{\prime}\log q\geq\sum_{q\leq x}\psi_{q}\log q.$
Hence, $\psi_{q}=\psi_{q}^{\prime}$ for all prime powers $q$. ∎
For an infinite extension $\mathcal{L}/\mathbb{Q}$, if $\psi_{q}>0$ for some prime power $q$, then over any tower =$\mathcal{L}=\{L_{i}\}$, all the primes above $p$ eventually almost split. More precisely, we have the following.
Proposition 4.3.
Let $\mathcal{L}/\mathbb{Q}$ be an infinite extension such that $\psi_{q}>0$ for some prime power $q=p^{k}$. For any $\epsilon>0$, there exists a number field $L\subset\mathcal{L}$ such that except for at most $\epsilon n_{L}$ prime ideals, all prime ideals in $L$ above $p$ almost split in $\mathcal{L}$.
Proof.
For a fixed prime $p$, by Lemma 4.1, we have
$\sum_{q=p^{k}}\psi_{q}\log q\leq\log p.$
Hence, for any $\epsilon>0$, there exists large $X$ such that
$\sum_{q=p^{k},q\geq X}\psi_{q}\log q<\epsilon/2.$
Let =$\mathcal{L}=\{L_{i}\}$ and say for all $i>N$, $\sum_{q=p^{k},q\geq X}\frac{\mathcal{N}_{q}(L_{i})}{n_{L_{i}}}\log q<\epsilon$. Again, by Lemma 4.1, we have
$\sum_{q\leq X,q=p^{k}}\frac{\mathcal{N}_{q}(L_{i})\,\log q}{[L_{i}:\mathbb{Q}]}\geq\sum_{q\leq X,q=p^{k}}\frac{\mathcal{N}_{q}(L_{i+1})\,\log q}{[L_{i+1}:\mathbb{Q}]}\geq\sum_{q\leq X,q=p^{k}}\psi_{q}\log q.$
Thus, for $i$ sufficiently large, we have
$\mathcal{N}_{q}(L_{i+1})=[L_{i+1}:L_{i}]\,\,\mathcal{N}_{q}(L_{i})+o(n_{L_{i}})$
for all $q\leq X$. Choosing $L=L_{i}$ for large enough $i$, all prime ideals in $L$ above $p$ with norm $\leq X$ almost split in $\mathcal{L}$. Furthermore, since $i>N$ the number of prime ideals in $L$ above $p$ with norm $>X$ is at most $\epsilon n_{L}$. This completes the proof. ∎
5. Proof of Theorem 1.1 and 1.3
Proof of Theorem 1.1.
Let $\mathcal{L}/\mathbb{Q}$ be an infinite extension. For a fixed prime $p$, let $\psi_{q}>0$ for some $q=p^{k}$. By Proposition 4.3, there exists a number field $L\subset\mathcal{L}$ such that all prime ideals in $L$ above $p$ with norm $\leq X$ almost split in $\mathcal{L}$. By Northcott’s theorem, there are finitely many $\alpha$ with degree $\leq n_{L}$ such that $h(\alpha)\leq\sum_{q}\psi_{q}\frac{\log q}{q+1}$, provided this sum converges.
For any $\alpha\in\mathcal{L}$, let $f_{L}(x)$ be its minimal polynomial over $L$ of degree $n$. Let $\mathfrak{p}\subset{\mathcal{O}}_{L}$ be a prime ideal above $p$ with norm $q\leq x$. Since $\mathfrak{p}$ is almost split in $L(\alpha)$, by Lemma 2.3, we can write
$f_{L}(x)=a_{n}(x-\alpha_{1})\,(x-\alpha_{2})\ldots(x-\alpha_{t})\,g(x),$
where $t=n+o(n)$. If $L_{\nu}$ is the completion of $L$ with respect to the prime ideal $\mathfrak{p}$, then we have $\alpha_{1},\alpha_{2},\cdots,\alpha_{t}\in L_{\nu}$. Hence, the valuation $\nu$ can be applied to all $\alpha_{i}$’s for $i\leq t$.
Let $P_{\omega}$ be the splitting field of $f_{L}(x)$ over $L_{\nu}$ with the unique valuation $\omega$ extending $\nu$ on $L_{\omega}$. Write
$f_{L}(x)=a_{n}(x-\beta_{1})(x-\beta_{2})\cdots(x-\beta_{n}),$
where $\beta_{i}\in P_{\omega}$ satisfy
$\omega(\beta_{1})\geq\cdots\geq\omega(\beta_{r})\geq 0>\omega(\beta_{r+1})\geq\cdots\geq\omega(\beta_{m}).$
The relative discriminant of $f_{L}$ is given by
$D(f_{L})$ $=a_{n}^{2n-2}\prod_{i<j}(\beta_{i}-\beta_{j})^{2}.$
Proceeding as in the proof of Corollary 1.2, we see that if at least one $\omega(\beta_{j})<0$, then
$\omega\left(\prod_{r+1}^{n}\prod_{i=1}^{j-1}(\beta_{j}-\beta_{i})\right)\geq\sum_{j=r+1}^{n}(j-1)\omega(\beta_{j}).$
Therefore
$\omega(D(f_{L}))$ $\geq 2\sum_{i<j\leq r}\omega(\beta_{j}-\beta_{i})-2\sum_{j=r+1}^{n}(n-j)\omega(\beta_{j})$
$\geq 2\sum_{\begin{subarray}{c}i<j\leq r\\ \beta_{i},\beta_{j}\in L_{\nu}\end{subarray}}\omega(\beta_{j}-\beta_{i})-2\sum_{\begin{subarray}{c}r<j\leq n\\ \beta_{j}\in L_{\nu}\end{subarray}}(n-j)\omega(\beta_{j}).$
Since all the $\beta_{j}$’s appearing in the RHS of the last inequality lie in $L_{\nu}$, we have
$\nu(D(f_{L}))\geq 2\sum_{\begin{subarray}{c}i<j\leq r\\ \beta_{i},\beta_{j}\in L_{\nu}\end{subarray}}\nu(\beta_{j}-\beta_{i})-2\sum_{\begin{subarray}{c}r<j\leq n\\ \beta_{j}\in L_{\nu}\end{subarray}}(n-j)\nu(\beta_{j}).$
Let $\mathbb{F}_{q}:=\mathcal{O}_{L}/\mathfrak{p}$ be the residue field of $\mathfrak{p}$. For each $x\in\mathbb{F}_{q}$, suppose $N_{x}$ denote the number of roots $\beta_{j}$ of $f_{L}$ in $L_{\nu}$ such that $\nu(\beta_{j}\geq 0)$, which lie in the residue class $x\bmod\mathfrak{p}$. Then, we have the lower bound
$\nu(D(f_{L})\geq 2\sum_{x\in\mathbb{F}_{q}}N_{x}(N_{x}-1)+2(n-r)(n-r-1).$
Since $(n-r)+\sum_{x\in\mathbb{F}_{q}}N_{x}=n+o(n)$, applying Cauchy-Schwarz inequality, we obtain
$\nu(D(f_{L}))$ $\geq\sum_{x\in\mathbb{F}_{q}}N_{x}(N_{x}-1)+(n-r)(n-r+1)$
$\geq\frac{n^{2}}{q+1}+o(n^{2}).$ (7)
where $q$ is the norm of $\mathfrak{p}$.
For any other rational prime power $q^{\prime}=p^{\prime k}$ such that $\psi_{q^{\prime}}>0$, the same analysis as above will yield a large number field $L^{\prime}\subset\mathcal{L}$ such that
$\nu^{\prime}(D(f_{L^{\prime}}))\geq\frac{n^{\prime 2}}{q^{\prime}+1}+o(n^{\prime 2}).$
Considering the compositum $LL^{\prime}\subset\mathcal{L}$, we obtain contributions from both the rational prime powers towards $D(f_{LL^{\prime}})$. Putting this together for all rational primes $\leq X$, there exists a large number field $\widetilde{L}\subset\mathcal{L}$ such that
$\log N_{\widetilde{L}/\mathbb{Q}}(D(f_{\widetilde{L}}))$ $\geq\sum_{\begin{subarray}{c}\nu\in M_{\widetilde{L}}\\ Norm(\nu)=q\leq X\end{subarray}}\nu(D(f))\log q$
$\geq\sum_{q\leq X}\mathcal{N}_{q}(L)\,\frac{\log q\,}{q+1}\,n^{2}+o(n^{2}).$
Using Lemma 3.1, we deduce
$\liminf_{\alpha\in\mathcal{L}}h(\alpha)\geq\sum_{q\leq X}\frac{1}{2}\psi_{q}\,\frac{\log q}{q}.$
Taking $X\to\infty$, we obtain Theorem 1.1. ∎
Proof of Theorem 1.3.
Recall that an infinite extension $\mathcal{L}/\mathbb{Q}$ is said be almost unramified at a prime $p$ if
$\sum_{\begin{subarray}{c}q=p^{k}\\ k\geq 1\end{subarray}}k\psi_{q}=1.$
In other words, if =$\mathcal{L}=\{L_{i}\}$,
$\sum_{k=1}^{\infty}k\,\mathcal{N}_{p^{k}}(L_{i})=n_{L_{i}}+o(n_{L_{i}}).$
We first note that an analog of Lemma 2.1 holds here, i.e., for $L\subset\mathcal{L}$, the number of prime ideals in $L$ with ramification index $\geq 2$ is $o(n_{L})$. Indeed, suppose $p\mathcal{O}_{L}=\mathfrak{p}_{1}^{e_{1}}\cdots\mathfrak{p}_{g}^{e_{g}}$. Since $\mathcal{L}$ is almost unramified, we have
$n_{L}=\sum_{q=p^{k}}k\,\mathcal{N}_{q}(L)+o(n_{L})=\sum_{j\leq g}f_{j}+o(n_{L}),$
where $f_{j}$ is the residue class index of $\mathfrak{p}_{j}$. Since $n_{L}=\sum_{j}e_{j}f_{j}$, we deduce that $e_{j}\geq 2$ for at most $o(n_{L})$ prime ideals above $p$.
Since $\mathcal{L}$ is almost unramified at $p$, given $\epsilon>0$ there exists an integer $N$ such that
$1-\epsilon\leq\sum_{\begin{subarray}{c}q=p^{k}\\ k\leq N\end{subarray}}k\psi_{q}\leq 1$
Thus for $\alpha\in\mathcal{L}$ and $L=\mathbb{Q}(\alpha)$ with $[L:\mathbb{Q}]$ large enough,
$(1-\epsilon)n_{L}\leq\sum_{k=1}^{N}k\,\mathcal{N}_{p^{k}}(L)\leq n_{L}.$
Let $f(x)=a_{n_{L}}x^{n_{L}}+a_{n_{L-1}}x^{n_{L-1}}+\cdots+a_{0}$ be the minimal polynomial of $\alpha$ over $\mathbb{Q}$ and $\widetilde{L}_{\omega}$ be the splitting field of $f(x)$ over $\mathbb{Q}_{p}$. Suppose the conjugates of $\alpha$ given by $\alpha_{1},\alpha_{2},\cdots,\alpha_{n_{L}}$ be ordered such that
$\omega({\alpha_{1}})\geq\omega({\alpha_{2}})\geq\ldots\geq\omega({\alpha_{r}})\geq 0>\omega({\alpha_{r+1}})\geq\ldots\geq\omega({\alpha_{n_{L}}}).$
Let $S_{1},S_{2},\ldots,S_{N}$ denote the set of unramified prime ideals in $L$ above $p$, with norm $p,p^{2},\ldots,p^{N}$ respectively. Thus, for each place $\nu_{k}$ in $S_{k}$, $L_{\nu_{k}}$ is an unramified extension of $\mathbb{Q}_{p}$ of degree $k$. By the uniqueness of the unramified extension of $\mathbb{Q}_{p}$ of a given degree and the fact that $\widetilde{L}_{\omega}$ is splitting field of $f(x)$, it follows that $L_{\nu_{k}}\subseteq\widetilde{L}_{\omega}$, for $1\leq k\leq N$, and each $L_{\nu_{k}}$ has exaclty $|S_{k}|$ many roots of $f(x)$ that generate $L_{\nu_{k}}$. Hence the roots of $f(x)\in\widetilde{L}_{\omega}$ that generate $L_{\nu_{k}}$ are in one to one correspondence with $S_{k}$.
For $1\leq k\leq N$, let $\alpha_{1,\,k},\,\alpha_{2,\,k},\ldots,\,\alpha_{|S_{i}|,\,k}$ be the roots of $f(x)$ that generate $L_{\nu_{k}}$. Grouping these roots together, we can write
[…]
$(x-\alpha_{1,\,N})\,(x-\alpha_{2,\,N})\,\ldots(x-\alpha_{|S_{N}|,\,N})\,.\,g(x)$
where $g(x)$ has degree $o(n_{L})$. The roots are ordered such that for a fixed $k$,
$\omega(\alpha_{1,\,k})\geq\omega(\alpha_{2,\,k})\geq\ldots\geq\omega(\alpha_{r_{k},\,k})\geq 0>\omega(\alpha_{r_{k}+1,\,k})\geq\omega(\alpha_{|S_{k}|,\,k}).$
Let $D(f)$ be the discriminant of $f(x)$. Then, analogous to (2), we have
$\omega(D(f))\,\,\,\geq$     $\,\,\,\,2\sum_{1\leq i<j\leq r}\omega(\alpha_{i}-\alpha_{j})-2\sum_{j=r+1}^{n_{L}}(n_{L}-j)\omega(\alpha_{j})$ (8)
$\geq$     $\,\,\,\,2\sum_{k=1}^{N}\left(\sum_{i<j\leq r_{k}}\omega(\alpha_{i,k}-\alpha_{j,k})-\sum_{j=r_{k}+1}^{|S_{k}|}(n_{L}-t_{j,k})\,\,w(\alpha_{j,k})\right),$
where $t_{j,k}$ is the integer such that $\alpha_{t_{j,k}}=\alpha_{j,k}$. The last inequality follows from the fact that each term on the RHS of (8) is non-negative and we are considering only those roots that generate $L_{\nu_{k}}$, for $1\leq k\leq N$.
Let $\nu$ denote the usual $p-$adic valuation on $\mathbb{Q}_{p}$. Since $w$ and $\nu_{k}$ are the unique extensions of $\nu$ to $\widetilde{L}_{\omega}$ and $L_{\nu_{k}}$ respectively, we have
$\omega(D(f))={\nu(D(f))}{e(\widetilde{L}_{\omega}|\mathbb{Q}_{p})},$
$\omega(\alpha_{i}-\alpha_{j})={\nu_{k}(\alpha_{i,l}-\alpha_{j,k})}{e(\widetilde{L}_{\omega}|L_{\nu_{k}})}\,\,\,\,\,\text{and}$
$\omega(\alpha_{j})={\nu_{k}(\alpha_{i})}{e(\widetilde{L}_{\omega}|L_{\nu_{k}})}.$
Furthermore, $L_{\nu_{k}}/\mathbb{Q}_{p}$ is unramified and hence $e(\widetilde{L}_{\omega}|L_{\nu_{k}})=e(\widetilde{L}_{\omega}|\mathbb{Q}_{p})$. Thus,
[…]
The roots $\alpha_{1,\,k},\,\alpha_{2,\,k},\ldots,\,\alpha_{r_{k},\,k}$ lie in $L_{\nu_{k}}$. Since the residue field of $L_{\nu_{k}}$ is $\mathbb{F}_{p^{k}}$, for each $x\in\mathbb{F}_{p^{k}}$, let $S_{x,k}$ denote the number of ${\alpha_{i,k}}$’s that is $x$ modulo $\mathbb{F}_{p^{k}}$. Also, note that the set $\{n_{L}-t_{j_{k}}\}$ appearing in (8) is a rearrangement of $\{n_{L}-j\}$ and hence is a set of distinct non-negative integers. Thus, we have the trivial bound
$-\sum_{j=r_{k}+1}^{|S_{k}|}(n_{L}-t_{j,k})\,\,\nu_{k}(\alpha_{j,k})\,\,\geq$ $\,\,\,\,\sum_{j=0}^{|S_{k}|-r_{k}-1}\,j\,\geq\,\,\frac{\left(|S_{k}|-r_{k}-1\right)\left(|S_{k}|-r_{k}\right)}{2}.$
Therefore
$\nu(D(f))\,\,\,\geq$ $\,\,\,\,\sum_{k=1}^{N}\left(\sum_{x\in\mathbb{F}_{p^{k}}}S_{x,k}\left(S_{x,k}-1\right)+\left(|S_{k}|-r_{k}-1\right)\left(|S_{k}|-r_{k}\right)\right)$
$\geq$ $\,\,\,\,\sum_{k=1}^{N}\left(\sum_{x\in\mathbb{F}_{p^{k}}}{S_{x,k}}^{2}+\left(|S_{k}|-r_{k}\right)^{2}\right)+O(n_{L}).$
Since $\left(|S_{k}|-r_{k}\right)+\sum_{x\in\mathbb{F}_{p^{k}}}S_{x,k}=|S_{k}|$, by Cauchy-Schwarz inequality, we obtain
$\nu_{k}(D(f))\,\,\,\geq$ $\,\,\,\,\sum_{k=1}^{N}\left(\frac{|S_{k}|^{2}}{p^{k}+1}\right)+O(n_{L}).$
As $|S_{k}|=l\,\mathcal{N}_{p^{k}}(L)$, we deduce
$\nu(D(f))\,\,\,\geq\,\,\,\,\sum_{k=1}^{N}\left(\frac{k^{2}{\mathcal{N}_{p^{k}}(L)}^{2}}{p^{k}+1}\right)+O(n_{L})\,\,\,\,\geq$ $\,\,\,\,{n_{L}}^{2}\,\,\sum_{k=1}^{N}\frac{k^{2}}{p^{k}+1}\left(\frac{\mathcal{N}_{p^{k}}(L)}{n_{L}}\right)^{2}\,+\,o({n_{L}}^{2}).$
Therefore,
$\log D(f)\,\,\geq\,\,{n_{L}}^{2}\,\,\sum_{k=1}^{N}\,\,\frac{k^{2}}{p^{k}+1}\left(\frac{\mathcal{N}_{p^{k}}(L)}{n_{L}}\right)^{2}\log q\,+\,o({n_{L}}^{2}).$
Now, using Theorem 2.2 and taking $N\rightarrow\infty$, we conclude
$\liminf_{\alpha\in\mathcal{L}}h(\alpha)\geq\frac{1}{2}\sum_{\begin{subarray}{c}q=p^{k}\\ k\geq 1\end{subarray}}k^{2}\,{\psi_{q}}^{2}\frac{\log q}{q+1}.$
6. Canonical heights on Elliptic Curves
Let $E$ be an elliptic curve defined over $\mathbb{Q}$ given by the equation $y^{2}=x^{3}+A\,x+B,$ where $A,\,B\in\mathbb{Q}$. For any extension $K/\mathbb{Q}$, denote by $E(K)$ the set of points $P\in E$ such that $x(P),\,y(P)\in K$. Define the Weil height of $P$ as
$h(P):=h(x(P)).$
Although Weil height is easy to define, it poses computational incompatibility with the group operation on elliptic curves. With this in mind, Néron and Tate introduced a new height function, namely the canonical height, defined as
$\widehat{h}(P)=\lim_{n\rightarrow\infty}\frac{h(2^{n}\,x(P))}{4^{n}},$
where $h$ is the usual Weil height. The canonical height $\widehat{h}$ is a quadratic form on $E$ and satisfies the following properties:
  1. (a)
    $\widehat{h}(P)\geq 0$ for all $P\in E$, and $\widehat{h}(P)=0$ iff $P$ is a torsion point.
  2. (b)
    $\widehat{h}([n]P)=n^{2}\,\widehat{h}(P)$ for all $n\in\mathbb{Z}$.
  3. (c)
    $\widehat{h}(A+B)+\widehat{h}(A-B)=2\widehat{h}(A)+2\widehat{h}(B)$.
In case of elliptic curves, the torsion points, having finite order can be thought of as analogs of roots of unity. Thus, property (a) is analogous to Kronecker’s theorem, which states that $h(\alpha)=0$ if and only if $\alpha$ is a root of unity. It is then natural to consider lower bounds of $\widehat{h}$ for non-torsion points on elliptic curves. In this aspect, Lehmer’s conjecture for elliptic curves, as formulated in [4] can be stated as follows.
Conjecture 6.1.
Let $E$ be an elliptic curve over a number field $K$. For any $P\in E(\overline{K})$, let $K(P)$ denote the field obtained by adjoining the coordinates of $P$ to $K$. Then, there exists a constant $C(E,K)$ depending only on $E$ and $K$ such that
$\widehat{h}(P)>\frac{C(E,K)}{[K(P):K]},$
for all non-torsion points $P\in E(\overline{K})$.
This remains open but weaker lower bounds have been obtained in special cases. For a fixed number field $K$ and $P\in E(\overline{K})$, denote by $D=[K(P):K]$. Some classical results are given below.
Table 1. Bounds on $\widehat{h}$ for $E(\overline{K})$
$\widehat{h}(P)\geqslant$ Restriction on $E$ Reference
$cD^{-10}(\log D)^{-6}$ none Anderson-Masser (1980) [4]
$cD^{-1}\left(\frac{\log\log(D)}{\log(D)}\right)^{3}$ $\mathrm{CM}$ Laurent (1983) [18]
$cD^{-3}(\log D)^{-2}$ none Masser (1989) [22]
$cD^{-2}(\log D)^{-2}$ $j$ non-integral Hindry-Silverman (1990) [15]
Analogous to the case of algebraic numbers, for an elliptic curve $E$ over $K$ and $S\subseteq\overline{K}$, $E(S)$ is said to have Bogomolov property (B) if there exists a constant $C(E,K)>0$, such that $\widehat{h}(P)>C(E,K)$, for all non-torsion points $P\in E(S)$.
Our objective is to obtain a lower bound on the canonical height of points on elliptic curves lying in certain infinite extensions of $\mathbb{Q}$. More precisely, suppose $\mathcal{K}/\mathbb{Q}$ is an infinite extension satisfying property (B) with respect to the Weil height. We would like to understand if $E(\mathcal{K})$ satisfies property (B) with respect to the canonical height. The first result of this kind is due to S. Zhang [34], who showed that for an elliptic curve $E/\mathbb{Q}$, the set of totally real points $E(\mathbb{Q}^{tr})$ satisfies property (B). This can be thought of as the analog of Schinzel’s result [26] for $\mathbb{Q}^{tr}$. In 2002, M. Baker proved property (B) for $E(K^{ab})$ when $E$ is non-CM or has non-integral $j$-invariant. This is the analogue of the theorem of Amoroso-Zannier [3] for $K^{ab}$. Moreover, Baker’s bound only depends on the $j-$invariant of $E$ and is effectively computable. Later, Silverman [30] proved this in full generality, without any condition on $E$, with the constant being ineffective.
Table 2. Bounds on $\widehat{h}$ for $E(K^{ab})$
$\widehat{h}(P)\geqslant$ Restriction on $E$ Reference
$cD^{-2}$ none Silverman (1981) [28]
$cD^{-1}(\log D)^{-2}$ none Masser (1989) [22]
$cD^{-2/3}$ $j$ nonintegral Hindry-Silverman (1990) [15]
$c$ $j$ nonintegral or CM Baker (2003) [5]
$c$ none Silverman (2003) [30]
We say a subfield $\mathcal{L}$ of $\overline{\mathbb{Q}}$ is totally $p$-adic of type $(e,f)$ if for any place $\nu$ in $\mathcal{L}$ above $p$, $\mathcal{L}_{\nu}/\mathbb{Q}_{p}$ is a finite extension with ramification index and residue class degree bounded above by $e$ and $f$ respectively. Let $E/K$ be an elliptic curve with semistable reduction at all places above $p\neq 2$. In 2005, M. Baker and C. Petsche [6] showed that if $\mathcal{L}/K$ is a totally $p$-adic field of type $(e,f)$ for a prime $p$, then $E(\mathcal{L})$ has property (B). This can be regarded as the elliptic analogue of equation (1). In this case, they also gave an effective upper bound for the number of torsion points in $E(\mathcal{L})$.
Let $E/\mathbb{Q}$ be an elliptic curve and $\mathbb{Q}(E_{tor})$ denote the field generated by all torsion points of $E$ in $\overline{\mathbb{Q}}$. In 2013, Habegger [14] showed that both $\mathbb{Q}(E_{tor})$ and $E(\mathbb{Q}(E_{tor}))$ have property (B) with respect to the Weil and the canonical height respectively. Note that the Kronecker-Weber theorem implies that $\mathbb{Q}^{ab}$ is generated by all the roots of unity in $\overline{\mathbb{Q}}$. Therefore, the field $\mathbb{Q}(E_{tor})$ can be thought of as the elliptic analogue of $\mathbb{Q}^{ab}$.
Let $\mathcal{L}/\mathbb{Q}$ be an asymptotically positive infinite extension. By Theorem 1.1, we know that $\mathcal{L}$ satisfies property (B). A natural question to consider is whether $E(\mathcal{L})$ satisfies property (B) with respect to the canonical height. We answer this in the affirmative below and prove it in the next section.
Theorem 6.2.
Let $E/\mathbb{Q}$ be an elliptic curve. Let $\mathcal{L}/\mathbb{Q}$ be an asymptotically positive infinite extension. Then
$\liminf_{P\in E(\mathcal{L})}\,\widehat{h}(P)\geq\sideset{}{{}^{\prime}}{\sum}_{\begin{subarray}{c}q=p^{k}\end{subarray}}\frac{1}{4}\psi_{q}(\mathcal{L})\frac{\log q}{q+1+2\sqrt{q}},$
where $\sideset{}{{}^{\prime}}{\sum}$ runs over all prime powers $q=p^{k}$ such that $E$ has good reduction at $p$. Therefore if $p$ is a prime of good reduction for $E$ and $\psi_{q}>0$ for $q=p^{k}$, then $E(\mathcal{L})$ has property (B).
7. Local heights and discrepancy
Let $K$ be a number field. Denote by $M_{K}$ and $M_{K}^{\infty}$ the set of all places and the set of archimedean places of $K$ respectively. Suppose that $\nu\in M_{K}\setminus M_{K}^{\infty}$ be a non-archimedean place corresponding to a prime ideal $\mathfrak{p}$ in $K$ above $p$. Denote by $|\cdot|_{\nu}$ the unique absolute value on $K_{\nu}$ which extends the usual $p$-adic absolute value on $\mathbb{Q}_{p}$.
Let $E$ be an elliptic curve over $K$ with identity $O$. Similar to Weil height, the canonical height $\widehat{h}$ on $E$ has a decomposition into local height functions $\lambda_{\nu}:E(K_{\nu})\rightarrow\mathbb{R}$ (called Néron local heights) for each place $\nu$ in $M_{K}$ (see [29, Chapter VI, Theorem 1.1]). For any $P\in E(K)\setminus\{O\}$, we have
$\widehat{h}(P)=\sum_{\nu\in M_{K}}\frac{[K_{\nu}:{\mathbb{Q}}_{\nu}]}{[K:\mathbb{Q}]}\lambda_{\nu}(P).$ (9)
Unlike the canonical height $\widehat{h}$, the local height functions $\lambda_{\nu}$ can take negative values, but at most at finitely many places. Let $\nu\in M_{K}$ be a finite place of good reduction for $E/K$. Choose a minimal Weierstrass equation for $E$ at $\nu$,
$y^{2}+a_{1}xy+a_{3}y=x^{3}+a_{2}x^{2}+a_{4}x+a_{6}.$
For a finite extension $L/K$, let $\omega\in M_{L}$ lie over $\nu$. Then by [30, Theorem 2.(c)]
$\widehat{\lambda}_{\omega}(P)=\max\{\omega(x/y),0\}\quad\text{ for all }P=(x,y)\in E(L)\setminus\{O\}.$
We recall some definitions and results from [6], where M. Baker and C. Petsche develop the theory of discrepancy on finite sets in elliptic curves. Let $E/\mathbb{Q}$ be an elliptic curve and $K$ be a number field. For a finite set $Z=\{P_{1},P_{2},\ldots,P_{N}\}\subseteq E(K)$, define
$\widehat{h}(Z)=\frac{1}{|Z|}\sum_{P\in Z}\widehat{h}(P)$
to be the average of canonical heights. The notion of discrepancy of a finite set $Z$ is a smooth version of $\frac{1}{N^{2}}\sum_{i,j}\widehat{h}(P_{i}-P_{j})$. For any non-archimedean place $\nu\in M_{K}$, the local discrepancy is defined as
$D_{\nu}(Z)=\frac{1}{N^{2}}\sum_{1\leq i,j\leq N}\lambda_{\nu}^{*}(P_{i}-P_{j}),$
where
$\lambda_{\nu}^{*}(P_{i}-P_{j})=\left\{\begin{array}[]{ll}\lambda_{\nu}(P_{i}-P_{j})&\mbox{if }P_{i}\neq P_{j}\\ j_{\nu}(P_{i},P_{j})&\mbox{if }P_{i}=P_{j}\end{array}\right.$
Here $j_{\nu}(P_{i},P_{j})$ is a bounded term (see [6, Section 3.1]). For an archimedean place $\nu\in M_{K}$, there is a one parameter family $\{\lambda_{t}\}_{t>0}$ of smooth functions $\lambda_{t}:E(\mathbb{C})\to\mathbb{R}$ such that $\lim_{t\to 0}\lambda_{t}=\lambda_{\nu}$ and $\frac{1}{N^{2}}\sum_{i,j}\lambda_{t}(P_{i}-P_{j})>0$ (see [17, section VI] and [11]). Now define the local discrepancy for archimedean places $\nu$ as
$D_{\nu}(Z)=\frac{1}{N^{2}}\sum_{1\leq i,j\leq N}\lambda_{1/N}(P_{i}-P_{j}).$
The global discrepancy for a finite set $Z\subset E(K)$, as introduced in [6] is defined as
$D(Z):=\sum_{\nu\in M_{K}}\frac{[K_{\nu}:\mathbb{Q}_{p}]}{[K:\mathbb{Q}]}D_{\nu}(Z),$
where $p$ is the corresponding rational prime such that $\nu|p$.
Proof of Theorem 6.2.
Our proof follows the methods in [6]. First note that since $\psi_{q}(\mathcal{L})>0$ for some $q=p^{k}$, there exists a finite extension of $\mathbb{Q}_{p}$, say $K_{\nu}$ such that there is an embedding $\mathcal{L}\subset K_{\nu}$. Since $|E_{tor}(K_{\nu})|<\infty$, we deduce that $|E_{tor}(\mathcal{L})|<\infty$. Now, we recall the height-discrepancy inequality [6, Theorem 4.1], which states that
$\mathcal{D}(Z)\leq 4\,\widehat{h}(Z)+\frac{1}{N}\left(\frac{1}{2}\log N+\frac{1}{12}h\left(j_{E}\right)+\frac{16}{5}\right),$ (10)
for any $Z=\{P_{1},P_{2},\cdots,P_{N}\}\subset E(K)$. For any $\nu\in M_{K}$ with good reduction at $E$ with norm $q=p^{k}$, we also have (see [6, Equation 6.36])
$D_{\nu}(Z)\geq\left(\frac{1}{m}-\frac{1}{N}\right)\log p,$
where $m=|E(\mathbb{F}_{q})|$ is the number of points of of $E$ in $\mathbb{F}_{q}$. By Hasse’s inequality,
$m\leq q+1+2\sqrt{q}$
and hence
$D_{\nu}(Z)\geq\left(\frac{1}{q+1+2\sqrt{q}}-\frac{1}{N}\right)\log p.$
Thus, the discrepancy
[…]
where $\sideset{}{{}^{\prime}}{\sum}$ runs over all $\nu\in M_{K}$ with good reduction over $E$ and $p$ the corresponding rational prime such that $\nu|p$ and $q=norm(\nu)$. Using the height-discrepancy inequality (10), we deduce
$4\widehat{h}(Z)\geq\sideset{}{{}^{\prime}}{\sum}_{q=p^{k}}\frac{\mathcal{N}_{q}(K)}{[K:\mathbb{Q}]}\left(\frac{1}{q+1+2\sqrt{q}}-\frac{1}{N}\right)\log q-\frac{1}{N}\left(\frac{1}{2}\log N+\frac{1}{12}h\left(j_{E}\right)+\frac{16}{5}\right).$
For any $N$, the above inequality holds for all possible $N$-points $\{P_{1},P_{2},\cdots P_{N}\}\subset E(K)$. Thus, for all but at most $N$-points in $E(K)$, we have $\widehat{h}(P)$ is greater than the RHS. Now, varying $K$ over $L_{i}$’s, where =$\mathcal{L}=\{L_{i}\}$ is a asymptotically positive extension of $\mathbb{Q}$ and taking $N$ large, we conclude that
$\liminf_{P\in E(\mathcal{L})}\,\widehat{h}(P)\geq\sideset{}{{}^{\prime}}{\sum}_{\begin{subarray}{c}q=p^{k}\end{subarray}}\frac{1}{4}\psi_{q}(\mathcal{L})\frac{\log q}{q+1+2\sqrt{q}},$
as required. ∎
8. Concluding remarks
The notion of relative height as discussed in Section 3 leads to several interesting questions. For instance, one can formulate a stronger Schinzel-Zassenhaus conjecture in this context. In [27], A. Schinzel and H. Zassenhaus proposed a weaker version of Lehmer’s conjecture, which was recently resolved by V. Dimitrov [9]. Let   $\alpha$  denote the house of $\alpha\in\overline{\mathbb{Q}}$ defined as
[…]
where $\alpha_{i}$’s are the conjugates of $\alpha$. Then, Dimitrov’s theorem states that for any non-zero $\alpha\in\overline{\mathbb{Q}}$ not a root of unity,
[…]
for $c=(\log 2)/4$. In the context of relative heights, it is possible to formulate a generalization of this conjecture in the following way. Let $m$ be a fixed positive integer. For any $\alpha\in\overline{\mathbb{Q}}$ define
[…]
Conjecture 8.1.
Let $m$ be a fixed positive integer. For all $\alpha\in\overline{\mathbb{Q}}\setminus\{0\}$, which are not roots of unity,
[…]
for an absolute constant $c>0$.
With regards to Theorem 6.2, it is possible to relax the condition that $E$ has good reduction at $p$ by using the potential theory developed in Fili-Petsche[12]. We relegate these investigations to future research.